50%
If you pick only one marble from the bag, at random, it can be any one of 26 marbles. Out of these, 5 of the marbles are green. Thus, there are 26 possible outcomes out of which 5 are favourable - to the event that the marble is green. Therefore the probability of a green marble is 5/26.
The calculations become more complicated if you consider choosing a green marble in several attempt: it depends on whether or not the marbles are replaced before the next one is picked.
What is the probability of getting a total of at least 9 in a single throw of two dice?
10/36. One of the easier ways to find an answer is simply to make a table showing the possibilities of every move.
1 2 3 4 5 6
1| 2 3 4 5 6 7
2| 3 4 5 6 7 8
3| 4 5 6 7 8 9
4| 5 6 7 8 9 10
5| 6 7 8 9 10 11
6| 7 8 9 10 11 12
Based on the table, there are 36 different roll combinations, with a total of 10 combinations that are at lest 9.
What is the difference between a probability mass function and a probability density function?
The probability mass function is used to characterize the distribution of discrete random variables, while the probability density function is used to characterize the distribution of absolutely continuous random variables.
You might want to read more about this at www.statlect.com/prbdst1.htm
(see the link below or on the right)
What is the probability of choosing a spade from a deck of cards?
Since there are 13 spades and 52 cards, the probability is 13/52 or 1/4=.25
What is the probability of drawing a diamond from a deck of 52 cards?
13/52 or 1/4 (25%)
Sample Space: Ace,2,3,4,5,6,7,8,9,10,Jack,Queen,King.
If you throw a 6 sided die 96 times how many times would it land on a factor of 24?
To solve this problem you: 1) Find all factors of 24 (there are 6 factors- 1,24,6,4,2,12) 2) You elemenate all numbers that are not on a six sided die (24,12) 3) Now use probability. The die is numbered 1-6 and you have numbers 1,2,4 and 6 which makes 4 out of 6. 4) This last step I can't show you, so I'll have to explain. You set up a proportion box. Write the 4 over the 6 on one side and next to the 6 write the 96. Over the 96 write N, which will represent the answer to your equation. Then multiply 4 times 96 and divide the product by 6. If you followed these steps you should get an answer of 64.
Find the probability of being dealt 5 diamonds from a standard deck of playing cards?
hypergeometric distribution: f(k;N,n,m) = f(5;52,13,5)
The probability of tails on the next toss is 1/2=.5 since this event is INDEPENDENT of the prior events.
If the two marbles are drawn without replacement, the probability is 16/33.
How many different arrangements of the letters in the word COLOR are there?
If you pick from each of the possible letters at each stage, then there are 5 possible options for the first letter. Then there are 4 possible options for the second. Then there are 3 possible options for the 3rd and so on. This leaves us with 5x4x3x2 = 120 different arrangements. However, each arrangement has been counted twice, as it doesn't matter which way around the Os go. So to get the true answer, we need to divide by 2. 120/2 = 60, and therefore the number of different arrangements of the letters in the word COLOR is 60.
We need to find two things. FIrst we need to know how many ways there are to roll a 3 and second we need to look at how many total events (the sample space) there are. So we could get a 3 with a 1 and 2 or a 2 and 1. Since we are rolling 2 dice we can't have 3 and 0 since we cant roll a 0. So there are 2 ways to get 3. Now Each die has 6 different outcomes when we roll it so using the multiplication law, we hae 36 possible outcomes. Of these 2 give us a sum of 3 so that is 2/36 or 1/18
Chances of rolling a 12 on a pair of dice?
The only way to get a 12 is to get two sixes. The chance of rolling a six is 1/6 and chance of rolling two sixes is 1/36 or about 2.78%
If A and B are mutually exclusive, P(A or B)=P(A) + P(B)
They both cannot occur together.
For example: A die is rolled. A = an odd number; B= number is divisible by 2.
P(A or B) = 1/3 + 1/3 = 2/3
If the probability of an event is 0.42 then what is the probability of its component?
I haven't heard of a component with regards to statistics. If, by chance, you are referring to the complement, it is the probability that the event does not occur. In this case, the complement would be 0.58.
What is the probability of picking a RED JACK from a deck of cards?
There are 2 red jacks, so 2/52 or about .038%
How many different arrangements of the letters of the word doodle are possible?
There are 180 different arrangements, you may punch in 6! to your calculator and get 720 and think this is right but you have to take into account that "d" and "o" and doubled, so, the arrangements will give you a different answer.
6*5*4*3*2*1/4= 180
180 arrangements. :)
This is a conditional probability, given the card is red, what is the chance it is a heart. Since there are 2 red hearts, the probability if 1/2
If a coin is flipped 7 times what is the probability of landing on all tails?
each time you flip the coin, probability to end on either side is 50% (or 0.5) (we disregard landing on the side).
So, to land on the same side 7 times, it is: 0.5^7
What is the probability of dying from jumping off a building?
This question cannot be answered.
The height of the building is what matters in this, and also how one hits the ground.
If you 'land' feet first, you might survive with serious injuries.
If you 'land' on your head you will most likely die immediately.
If the building is anything over 25 feet you will most likely die.
---
Jumping off is not the problem - it's the sudden stop that does the damage. Base-jumpers 'fall' off buildings regularly and live, even from enormous heights - they have parachutes. And stuntmen fall off buildings and out high windows and live because they either land on piled up cardboard boxes or into a swimming pool of have some other way of cushioning their impact.
What is probability of getting all heads if 3 coins are tossed simultaneously?
There is a probability of 0.5 of heads on each coin, so there is a 1/2 * 1/2 *1/2 = 1/8 = 0.125 chance that all coins land heads. You can also express this as a percentage, 12.5% chance. The odds are 1 in 8 that this can happen.